OpenStudy (anonymous):

Solve the equation 2a-9=2(a+1) +13a+4

OpenStudy (abmon98):

collect a terms together and expand the bracket 2a-9=2a+2+13a+4 2a-9=15a+6 13a=-15 a=-15/13

OpenStudy (anonymous):

would you mind showing me how you got that

OpenStudy (abmon98):

we expand the brackets by multiply the a by 2 and 1 by 2 also. 2a-9=2a+2 +13a+4 now we are trying to solve to obtain the value of a so we collect all the a-term on a side and integers on the other side -9-2-4=-2a+2a+13a -9-6=13a -15=13a we divide both sides by 13 -15/13=13a *1/13 -15/13=a